# Thunderdome: Timelock-Free Rationally-Secure Virtual Channels

### v3

@@ Appendix A Byzantine Model Security Proof
For the first case, assume Alice initiate unilateral closing first. After receiving $12f+$ 1 signatures from the wardens, Alice’s contract will send the state notification to Bob’s contract. According to Protocol 5, since Bob’s contract has not yet received enough signatures from its wardens, it will store the state received from Alice’s contract and reply with a Null. Subsequently, regardless of the state later published to Bob’s contract, it will only close with the state received from Alice’s contract. Conversely, if Bob initiates the closing first, then both parts of Thunderdome will close with Bob’s state. In the event that Alice and Bob initiate unilateral closings simultaneously, both contracts will receive two different states—one from Alice and one from Bob. By comparing the sequence numbers, both contracts will decide on Bob’s state, as it has the higher sequence number. Thus, the first case cannot occur.
@@ Appendix C Unilateral Closing Game
$PF$ Number of valid proofs-of-fraud she publishes on-chain
@@
[…]
During the unilateral closing process, multiple wardens participate in the game. For simplicity, we assume that each payment channel committee consists of $103f+1=$ 10 wardens. According to Protocol 4, at least 7 of these wardens from the closing party’s warden committee are required to publish states onchain. To further simplify the analysis, we divide these 10 wardens into three groups, denoted as $\{W_{1},W_{2},W_{3}\}$. Specifically, $W_{1}$ contains 3 wardens who are unaware of the latest virtual channel state, while $W_{2}$ and $W_{3}$ consist of 4 and 3 wardens, respectively, who know the latest state. Since all wardens are treated identically in this analysis, we focus only on the total number of wardens adopting a particular strategy. Additionally, we introduce the following assumption for wardens’ strategies:
@@ Appendix D Game-theoretic Analysis (omitted proofs)
Assume the current state of both the Alice-Ingrid and Ingrid-Bob payment channels is $(a,b)$. If all three parties agree to open a virtual channel between Alice and Bob with an initial state of $(c,d)$, both channels are updated to $(a-c,b-d)$. A security concern arises if the funding transactions differ. For example, Ingrid could collude with Alice and mislead Bob by updating the Ingrid-Bob channel with an amount $d^{\prime}\neq d$. If $d^{\prime}>d$, Ingrid may incur a loss, as Alice would be able to pay Bob more coins in the virtual channel than she has locked in the Alice-Ingrid channel. Conversely, if $d^{\prime}<d$, Alice could lose funds. Since either Alice or Ingrid would face a financial loss in this scenario—and both are modeled as utility-maximizing agents—such an incorrect opening would not occur. ∎
@@
As shown in Fig. 11, the final utility for wardens depends on the probability $p\in\{p_{1},p_{2}|p_{1}+p_{2}=1\}$ that Ingrid has knowledge of the Thunderdome state. We simplify the original unilateral closing game $\mathcal{G}_{I}$ by categorizing the warden’s strategy into two types. The honest strategy, $\sigma^{h}_{W}(\bot)=\{3P_{l},4P_{l},/\}$, represents each warden publishing the latest state they know. The dishonest strategy, $\sigma^{dish}_{W}(\bot)=\{3P_{l},aP_{l}+bP_{o},/\}$, represents $b$ wardens publishing outdated states on-chain, where $a+b=4$ and $40\leq a<$ 4. The warden’s expected utility for the dishonest strategy $\sigma^{dish}W(\bot)$ is $u_{dish}=p_{1}(ak-bc)+4p_{2}k=(ap_{1}+4p_{2})k-bcp_{1}$. The expected utility for the honest strategy $\sigma^{h}_{W}(\bot)$ is $u_{h}=4k$. Given that $a<4$ and $p_{1}>0$, it follows that $u_{dish}=(ap_{1}+4p_{2})k-bcp_{1}<4k=u_{h}$. Therefore, the honest strategy $\sigma^{h}_{W}(\bot)=\{3P_{l},4P_{l},/\}$ yields a higher utility, proving that Lemma 2 still holds for the closing party without knowledge of the Thunderdome state. ∎
@@
In the game-theoretic model, wardens could be bribed to sign two different states during a Thunderdome update. However, since we require $12f+$ 1 accountable wardens to sign the state—wardens who can be penalized for such misbehavior—the security of both the update and unilateral closing processes can be analyzed using the same game model. This model discussed in Appendix C, yields the same result for the update protocol. The proof for Lemma 2 also applies to the update protocol. In this context, $P_{l}$ represents honestly signing a single state with the latest sequence number, while $P_{o}$ indicates signing multiple states with the same sequence number. During the closing procedure, the closing party can still punish double-signing through fraud proofs, just as outdated state publishing is penalized. Therefore, for the update protocol, the SPNE strategy for wardens remains $P_{l}$, meaning they sign only one state. Additionlly, the proof for Lemma 3 also applies to the update protocol and implies that wardens cannot determine whether Ingrid is able to punish them for signing two different states. ∎
@@
$Ignore$ $Agree$ $Disagree$
$Uni$ $(\alpha-\epsilon,\alpha)$ $(\alpha-\epsilon,\alpha)$ $(\alpha-\epsilon,\alpha)$
$Old$ $(\alpha-\epsilon,\alpha)$ $(\alpha+d_{B},\alpha-d_{B})$ $(\alpha,\alpha-\epsilon)$
$New$ $(\alpha-\epsilon,\alpha)$ $(\alpha,\alpha)$ $(\alpha,\alpha-\epsilon)$
@@ E.2 Security under game-theoretic model
To show the security of the closing game, we first prove that the unilateral closing subgame is secure. The update protocol and the unilateral closing game remain the same as in the two-hop case, requiring $12f+$ 1 wardens from the local committee to publish the state. Additionally, wardens cannot determine whether intermediary parties (e.g., Bob and Carol) know the Thunderdome state, as collusion between the intermediary and the main Thunderdome parties (like Alice and Dave) is always possible – similar to the two-hop Thunderdome scenario. Therefore, the security of the unilateral closing subgame is inherited from the security analysis of the two-hop Thunderdome presented earlier.

### v4

@@ Appendix A Byzantine Model Security Proof
For the first case, assume Alice initiate unilateral closing first. After receiving $2f+1$ signatures from the wardens, Alice’s contract will send the state notification to Bob’s contract. According to Protocol 5, since Bob’s contract has not yet received enough signatures from its wardens, it will store the state received from Alice’s contract and reply with a Null. Subsequently, regardless of the state later published to Bob’s contract, it will only close with the state received from Alice’s contract. Conversely, if Bob initiates the closing first, then both parts of Thunderdome will close with Bob’s state. In the event that Alice and Bob initiate unilateral closings simultaneously, both contracts will receive two different states—one from Alice and one from Bob. By comparing the sequence numbers, both contracts will decide on Bob’s state, as it has the higher sequence number. Thus, the first case cannot occur.
@@ Appendix C Unilateral Closing Game
PFPF Number of valid proofs-of-fraud she publishes on-chain
@@
$P_{l}$ Publish the honest state $\backslash$ Sign honestly
$P_{o}$ Publish the dishonest state $\backslash$ Sign dishonestly
During the unilateral closing process, multiple wardens participate in the game. For simplicity, we assume that each payment channel committee consists of $3f+1=10$ wardens. According to Protocol 4, at least 7 of these wardens from the closing party’s warden committee are required to publish states onchain. To further simplify the analysis, we divide these 10 wardens into three groups, denoted as $\{W_{1},W_{2},W_{3}\}$. Specifically, $W_{1}$ contains 3 wardens who are unaware of the latest virtual channel state, while $W_{2}$ and $W_{3}$ consist of 4 and 3 wardens, respectively, who know the latest state. Since all wardens are treated identically in this analysis, we focus only on the total number of wardens adopting a particular strategy. Additionally, we introduce the following assumption for wardens’ strategies:
@@ Appendix D Game-theoretic Analysis (omitted proofs)
Assume the current state of both the Alice-Ingrid and Ingrid-Bob payment channels is $(a,b)$. If all three parties agree to open a virtual channel between Alice and Bob with an initial state of $(c,d)$, both channels are updated to (ac,bd)(a-c,b-d). A security concern arises if the funding transactions differ. For example, Ingrid could collude with Alice and mislead Bob by updating the Ingrid-Bob channel with an amount $d^{\prime}\neq d$. If $d^{\prime}>d$, Ingrid may incur a loss, as Alice would be able to pay Bob more coins in the virtual channel than she has locked in the Alice-Ingrid channel. Conversely, if $d^{\prime}<d$, Alice could lose funds. Since either Alice or Ingrid would face a financial loss in this scenario—and both are modeled as utility-maximizing agents—such an incorrect opening would not occur. ∎
@@
As shown in Fig. 11, the final utility for wardens depends on the probability $p\in\{p_{1},p_{2}|p_{1}+p_{2}=1\}$ that Ingrid has knowledge of the Thunderdome state. We simplify the original unilateral closing game $\mathcal{G}_{I}$ by categorizing the warden’s strategy into two types. The honest strategy, $\sigma^{h}_{W}(\bot)=\{3P_{l},4P_{l},/\}$, represents each warden publishing the latest state they know. The dishonest strategy, $\sigma^{dish}_{W}(\bot)=\{3P_{l},aP_{l}+bP_{o},/\}$, represents $b$ wardens publishing outdated states on-chain, where $a+b=4$ and $0\leq a<4$. The warden’s expected utility for the dishonest strategy $\sigma^{dish}W(\bot)$ is $u_{dish}=p_{1}(ak-bc)+4p_{2}k=(ap_{1}+4p_{2})k-bcp_{1}$. The expected utility for the honest strategy $\sigma^{h}_{W}(\bot)$ is $u_{h}=4k$. Given that $a<4$ and $p_{1}>0$, it follows that $u_{dish}=(ap_{1}+4p_{2})k-bcp_{1}<4k=u_{h}$. Therefore, the honest strategy $\sigma^{h}_{W}(\bot)=\{3P_{l},4P_{l},/\}$ yields a higher utility, proving that Lemma 2 still holds for the closing party without knowledge of the Thunderdome state. ∎
@@
In the game-theoretic model, wardens could be bribed to sign two different states during a Thunderdome update. However, since we require $2f+1$ accountable wardens to sign the state—wardens who can be penalized for such misbehavior—the security of both the update and unilateral closing processes can be analyzed using the same game model. This model discussed in Appendix C, yields the same result for the update protocol. The proof for Lemma 2 also applies to the update protocol. In this context, $P_{l}$ represents honestly signing a single state with the latest sequence number, while $P_{o}$ indicates signing multiple states with the same sequence number. During the closing procedure, the closing party can still punish double-signing through fraud proofs, just as outdated state publishing is penalized. Therefore, for the update protocol, the SPNE strategy for wardens remains $P_{l}$, meaning they sign only one state. Additionlly, the proof for Lemma 3 also applies to the update protocol and implies that wardens cannot determine whether Ingrid is able to punish them for signing two different states. ∎
@@
IgnoreIgnore AgreeAgree DisagreeDisagree
UniUni $(\alpha-\epsilon,\alpha)$ $(\alpha-\epsilon,\alpha)$ $(\alpha-\epsilon,\alpha)$
OldOld $(\alpha-\epsilon,\alpha)$ $(\alpha+d_{B},\alpha-d_{B})$ $(\alpha,\alpha-\epsilon)$
NewNew $(\alpha-\epsilon,\alpha)$ $(\alpha,\alpha)$ $(\alpha,\alpha-\epsilon)$
@@ E.2 Security under game-theoretic model
To show the security of the closing game, we first prove that the unilateral closing subgame is secure. The update protocol and the unilateral closing game remain the same as in the two-hop case, requiring $2f+1$ wardens from the local committee to publish the state. Additionally, wardens cannot determine whether intermediary parties (e.g., Bob and Carol) know the Thunderdome state, as collusion between the intermediary and the main Thunderdome parties (like Alice and Dave) is always possible – similar to the two-hop Thunderdome scenario. Therefore, the security of the unilateral closing subgame is inherited from the security analysis of the two-hop Thunderdome presented earlier.